Question 2: draw a problem that needs the slope of a line formula and demonstrate
and give examples with different missing
measurement we need to find to slope of a line formula clearly understandable
Answer #1

① Example : 3 + 2 4 5 G line det of the line is given mentioned above, from the We have two points. (2,4)= (6,3) (x2, Ya )= (Example : 3 4 S. we from the given line get two points let (441) = (3,-1) (xo. 92) =(4,-2) slope of the tine Ya-Y, m = -> m=